OpenStudy (anonymous):

Ok. I'd really appreciate the answers to these questions. I'm in a hurry & I need to get this finished. 1)-2 <= d + 1/2 < 4 1/2 2)2t <= -4 or 7t >= 49

OpenStudy (anonymous):

-2 <= d + 1/2 < 4 1/2 -2-1/2<= d < 4 1/2-1/2 -2 1/2<= d < 4

OpenStudy (anonymous):

2)2t <= -4 or 7t >= 49 /2 /2 /7 /7 t <= -2 or t >= 7

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

dont forget the medal :p

OpenStudy (anonymous):

right right. xD